# A Culturally-diverse Multilingual Multimodal Video Benchmark & Model

### v1

@@ 7 Limitations
ViMUL, to the best of our knowledge, is the first multilingual Video LMM benchmark that exhibits cultural and linguistic inclusivity across 14 languages and 15 diverse domains. However, despite being the first benchmark for evaluating Video LMMs in a multilingual setting, it has certain limitations.
First, the sub-categories in ViMUL-Bench contain more general VQA samples compared to those in the cultural section. This imbalance is primarily due to the high cost of curating culturally grounded data, which requires extensive human effort. Furthermore, ViMUL-Bench includes more short and medium-length videos than long ones. High-quality long videos are particularly difficult to source for low-resource languages such as Sinhala, Tamil, and Urdu, and require additional human verification to generate reliable QA pairs. Although we have conducted cycle consistency checks on the multilingual instruction set fine-tuning, further human validation could enhance the quality of the results. Lastly, while the ViMUL model serves as a simple baseline for multilingual video LMMs, it highlights the performance trade-off between high- and low-resource languages in video understanding.
As future research directions, our work can be extended to include additional languages, particularly those with limited digital representation. Furthermore, developing culturally-specific, large-scale training datasets tailored explicitly to underrepresented communities would further enhance the inclusivity and effectiveness of multilingual video LMMs. We also expect culture-specific video-language preference data collection can help improve LMMs’ performance further via reinforcement learning.
@@ 8 Ethical Consideration
Our work reports a standardized multilingual video LMM evaluation benchmark. We hope both ViMUL and ViMUL-Bench will contribute to more consistent evaluation across diverse domains, particularly for under-represented languages in VidLMM research. Since the cultural videos in ViMUL-Bench are sourced from Internet, some domains may be under-represented, leading to potential biases. To ensure high translation quality, fluent native speakers thoroughly reviewed and verified our initial GPT-4o-generated translations for consistency and accuracy. The verification involves 16 volunteers from diverse linguistic backgrounds for benchmark curation. Annotators were required to have familiarity with the cultural context of the specific country-language pair they worked on. Additional demographic details about annotators are presented in Sec. D suppl. material.
@@ Appendix D Volunteer Demographics
Figure 15: The top left figure shows the percentage of our volunteers with respect to linguistic skill level. The top right shows thir age distribution. Then in bottom left we have their active years of communication, followed by the duration of their stay in their respective countries.
We have a total of 16 volunteers form various backgrounds who assisted us in curating and verifying our ViMUL Bench. Around 77.8% of our verifiers are native or bilingual, followed by 11.1% of them proficient and the rest are intermediate in their respective languages. Around 94.4% stayed for more than 15 years in their country, where they learned their first language. We can observe in 15 our contributors range from the age bracket of 18 - 25, 25 - 30 and onwards, which makes the age distribution more diverse, and includes prior experience. In terms of geographical distribution, they come from the following countries as follows Bangladesh, China, Germany, India, Japan, Lebanon, Morocco, Pakistan, Sri Lanka, Sweden, Ukraine and USA. This diversity in geo-location helps us to cater to cultural nuances of that corresponding region, hence allowing us to get translations and verifications which are authentic to that place.
@@ Appendix F Performance of Phi-4 as a judge
Figure 17: Performance of Phi-4 scores compared to human score. The figure illustrates that Phi-4 scores are similar to human scores where they differ most in Japanese by 4.1% and the least in Chinese by 0.4%
@@ Appendix I Performance based upon frames
Appendix N Experimental Setup and Hyperparameters

### v2

@@ 6 Conclusion
Acknowledgement: The computations were enabled through resources provided by NAISS at Alvis, partially funded by the Swedish Research Council under grant agreement no. 2022-06725, LUMI, hosted by CSC (Finland) and the LUMI consortium, and the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC. This work was partially supported by the Swedish Research Council (2022-04266), from KAW (DarkTree project; 2024.0076) and VR starting grant (2016-05543).
@@ 7 Limitations
ViMUL, to the best of our knowledge, is the first multilingual Video LMM benchmark that exhibits cultural and linguistic inclusivity across 14 languages and 15 diverse domains. However, despite being the first benchmark for evaluating Video LMMs in a multilingual setting, it has certain limitations. ViMUL-Bench contains more general VQA samples in its sub-categories than in the cultural section, largely due to the high cost of curating culturally grounded data, which demands extensive human effort. It also has more short and medium-length videos than long ones, as high-quality long videos are harder to source for low-resource languages such as Sinhala, Tamil, and Urdu, and require extra verification for reliable QA pairs. While we applied cycle consistency checks during multilingual instruction fine-tuning, further human validation could improve quality. Finally, the ViMUL model, though a simple baseline, highlights the trade-off between high- and low-resource languages in video understanding. As future research directions, our work can be extended to include additional languages, particularly those with limited digital representation. Furthermore, developing culturally-specific, large-scale training datasets tailored explicitly to underrepresented communities would further enhance the inclusivity and effectiveness of multilingual video LMMs. We also expect culture-specific video-language preference data collection can help improve LMMs’ performance further via RL.
@@ 8 Ethical Consideration
Our work reports a standardized multilingual video LMM evaluation benchmark. We hope both ViMUL and ViMUL-Bench will contribute to more consistent evaluation across diverse domains, particularly for underrepresented languages in VidLMM research. Since the cultural videos in ViMUL-Bench are sourced from Internet, some domains may be under-represented, leading to potential biases. To ensure high translation quality, fluent native speakers thoroughly reviewed and verified our initial GPT-4o-generated translations for consistency and accuracy. The verification involves 16 volunteers from diverse linguistic backgrounds, requiring having familiarity with the cultural context of the specific country-language pair they worked on. Additional annotator demographic details are presented in Sec. D suppl. material.
@@ Appendix D Volunteer Demographics
Figure 15: The top left figure shows the percentage of our volunteers with respect to linguistic skill level. The top right shows their age distribution. Then in bottom left we have their active years of communication, followed by the duration of their stay in their respective countries.
We have a total of 16 volunteers from various backgrounds who assisted us in curating and verifying our ViMUL Bench. Among these, two volunteers are proficient in two different languages each, and they contributed separately for both languages. To accurately reflect their contributions and language proficiency levels, we consider a total of 18 language instances rather than just 16 individuals when calculating statistics.
Around 77.8% of our verifiers are native or bilingual, followed by 11.1% of them proficient and the rest are intermediate in their respective languages. Around 94.4% stayed for more than 15 years in their country, where they learned their first language. We can observe in Fig. 15 our contributors range from the age bracket of 18 - 25, 25 - 30 and onwards, which makes the age distribution more diverse, and includes prior experience. In terms of geographical distribution, they come from the following countries as follows Bangladesh, China, Germany, India, Japan, Lebanon, Morocco, Pakistan, Sri Lanka, Sweden, Ukraine and USA. This diversity in geo-location helps us to cater to cultural nuances of that corresponding region, hence allowing us to get translations and verifications which are authentic to that place.
@@ Appendix F Performance of Phi-4 as a judge
Figure 17: Performance of Phi-4 scores compared to human score. The figure illustrates that Phi-4 scores are similar to human scores where they differ most in Japanese by 4.1% and the least in Chinese by 0.4%.
@@ Appendix I Performance based upon frames
Appendix N License for Artifacts
We use the following datasets and models in our paper. We list all the licenses for each of them below
  • Base LLMs and the multimodal finetuned models accessed via Hugging Face (e.g., Video-ChatGPT-7B, LLaVA-NeXT-Video-7B,, VideoChat2-Mistral-7B, VILA v1.5 - 7B, LLaVA OneVision - 7B, GPT-4o, GPT-4o-mini, Qwen-3, LLaMA-3.1, Phi-4) are cited with their respective papers, model cards, and URLs.
    Licenses for multimodal LMMs:
    • Video-ChatGPT-7B: CC BY 4.0
    • LLaVA-NeXT-Video-7B: LLAMA 2 Community License
    • VideoChat2-Mistral-7B: Apache License 2.0
    • VILA-v1.5-7B: Apache License 2.0 (code), CC BY-NC-SA 4.0 (weights)
    • LLaVA-OneVision-7B: LLAMA 2 Community License
    • Qwen-3: Apache License 2.0
    • Phi-4: MIT License
    • GPT-4o: Closed-source (OpenAI Proprietary)
    • GPT-4o-mini: Closed-source (OpenAI Proprietary)
    • LLaMA-3.1: LLaMA 3 Community License (restricted commercial use)
    Licenses for the datasets used in our work for the construction of fine-tuning and evaluation benchmark.
    • CVRR: CC BY-NC-SA 4.0
    • VideoMME: Academic research only; commercial use prohibited
    • VCG Diverse: MIT License
    • MVBench: MIT License
    • Video-Instruct100K: CC BY-SA 4.0
    • LLaVA-Video-178K: Apache License 2.0
    • NeXT-QA: MIT License
    • PerceptionTest: CC BY 4.0
    • CLEVRER: CC0 License
    • ActivityNetQA: MIT License (non-commercial use only)
  • The video samples for cultural category was scraped manually from the publicly accessible websites, such as YouTube. The usage of this content is compliant with fair-dealing law for non-commercial academic research. We do not redistribute the original video and text data under commercial licensing.
  • Our codebase builds on open-source frameworks such as PyTorch (BSD-style license) and Transformers (Apache 2.0). Our evaluation framework is based on the lmms-eval framework, based on Apache License 2.0.
The terms of use for each resource are respected, and we provide links and citations in the supplementary material and code repository. We release our work under the license CC BY-NC 4.0.
Appendix O Documentation of Artifacts
We list the coverage of domains of the datasets used in our paper:
O.1 Evaluation Benchmark Domains
  • CVRR:
    • Non-existent actions with non-existent scene depictions
    • Non-existent actions with existent scene depictions
    • Time order understanding
    • Understanding of emotional context
    • Interpretation of social context
    • Unusual and Physically Anonymous activities
    • Continuity and Object Instance Count
    • Interpretation of visual context
    • Partial actions
    • Fine-grained action understanding
    • Multiple actions in a single video
  • VideoMME:
    • Domains in dataset:
        Life Record: Travel, Daily Life, Fashion, Food, Handicraft, Pet & Animal, Exercise, Multilingual
        Artistic Performance: Acrobatics, Variety Show, Magic Show, Stage Play
        Sports Competition: Other Sports, Athletics, Football, Basketball, Esports
        Film & Television: News Report, Documentary, Movie & TV Show, Animation
        Knowledge: Technology, Life Tip, Law, Geography, Astronomy, Finance & Commerce, Biology & Medicine, Literature & Art, Humanity & History
    • Domains in various questions:
        Temporal Perception
        Spatial Perception
        Attribute Perception
        Action Recognition
        Object Recognition
        OCR Problems
        Counting Problems
        Temporal Reasoning
        Spatial Reasoning
        Action Reasoning
        Object Reasoning
        Information Synopsis
  • MVBench:
    • Spatial Understanding
        Action: What’s the man doing?
        Object: What’s on the table?
        Position: Is the man on the stage?
        Count: How many chairs?
        Scene: Where’s the man?
        Pose: What’s the man’s pose?
        Attribute: What color is the desk?
        Character: What are the subtitles?
        Cognition: Why is the man singing in the canteen?
    • Temporal Understanding
        Action: Action Sequence, Action Antonym, Action Prediction, Unexpected Action, Fine-grained Action
        Object: Object Shuffle, Object Existence, Object Interaction
        Position: Moving Direction, Action Localization
        Count: Action Count, Moving Count
        Scene: Scene Transition
        Pose: Fine-grained Pose
        Attribute: State Change, Moving Attribute
        Character: Character Order
        Cognition: Episodic Reasoning, Egocentric Navigation, Counterfactual Inference
  • VCG-Diverse
    • Video Domains
        News, Surveillance, Traffic, Automobiles, Sports, Gaming, Cooking, HowTo, Travel, Pets, Education, Science, Entertainment, Music, Comedy, Film, Lifestyle, Activism
    • Question Types
        Sequential Understanding: Cooking, How-to, Education
        Predictive Reasoning: Sports, Gaming
        World Knowledge: Science, News
        Causal Reasoning: Surveillance, Activism
        Emotional Reasoning: Entertainment, Film, Comedy
        Analytical Reasoning: Traffic, Automobile
O.2 Instruction Fine-Tuning Domains
  • Video-Instruct100K: Video Summarization, Description-based QA (spatial, temporal, relational, reasoning), Creative/Generative QA
  • LLaVA-Video-178K: Academic Sources, YouTube Videos, ActivityNetQA, NeXT-QA, PerceptionTest, LLaVA-Hound
  • NeXT-QA: Causal Action Reasoning, Temporal Action Reasoning, Object Interaction Understanding, Daily Activity Scenarios
  • PerceptionTest: Memory, Abstraction, Physics, Semantics
  • CLEVRER: Temporal Reasoning, Causal Reasoning, Physical Dynamics, Symbolic Event Representation
  • ActivityNetQA: Long-Term Spatio-Temporal Reasoning, Complex Web Video Understanding, Human-Annotated QA Pairs
Appendix P Model Size and Budget
We had the following compute budget for our project:
  • Evaluation on ViMUL-Bench (including ablations):
    • Total GPU Hours: 9
    • GPU Variant: AMD-MI200 (64GB)
  • Finetuning ViMUL on ViMUL-Instruct:
    • Total GPU Hours: 200–240
    • GPU Variant: AMD-MI200 (64GB)
  • Translation of 130k samples for cycle consistency using Qwen-3:
    • Total GPU Hours: 12,000
    • GPU Variant: AMD-MI200 (64GB)
Appendix Q Experimental Setup and Hyperparameters
@@
Appendix R Instruction to Volunteers
R.1 Cultural Video Dataset Curation
For each language, we have identified eight distinct cultural categories. We gave the following instructions to the volunteers to collect the videos:
  • From your selected language choose 3 to 4 videos from each of the 8 cultural categories
  • Please avoid choosing videos that contain sensitive personal information such as private addresses, Social Security numbers, or any other confidential data.
  • Whenever possible, select videos filmed in public places to respect privacy. Ensure that the video content does not infringe on the privacy of any individuals or groups.
  • Videos should not depict or disclose any private or sensitive content without the consent of the people involved.
  • Always ensure that the content of the video respect the privacy of individuals and do not include private or sensitive information.
  • If there is any doubt about the appropriateness of a video, please consult the project supervisor or team for clarification.
  • Ensure that the video is public and its license is also public.
R.2 Cultural QA Curation
After the videos are collected, we asked our volunteers to curate QA pairs of both multiple choice and open-ended QA pairs, for each video in their respective native language. The instructions were given as follows:
  • Please watch each video carefully. If you feel that the video does not align with the specified category or the native language, you are welcome to replace it with a new YouTube video link. Ensure that the video is public and its license is also public.
  • After watching the video, you are required to create 3 Multiple Choice Questions (MCQs) and 1 Short Answer Question (SAQ) in English based on the video’s content.
  • Additionally, for each question, you must provide a translated version in the native language along with the answer.
Instructions for Writing Multiple Choice Questions (MCQs)
  • Video-Dependent Questions: Ensure that each question requires the viewer to watch the video to answer. Avoid questions that could be answered with general knowledge or basic text comprehension. For example: "Where is the University of Central Florida?" is incorrect, while "Where is the university shown in the video?" is correct.
  • Contextual Understanding: The question should assess the viewer’s understanding of the video’s context, including spatial relationships, object interactions, and scene transitions. For example: "Where does the car park after crossing the street, as shown in the video?" is a suitable question.
  • Avoid Ambiguity: Ensure that all questions and answer options are clear, unambiguous, and directly linked to the video content. Vague or open-ended questions can confuse models and reduce the benchmark’s effectiveness. For example: Instead of asking "What happens next?", use "What does the person do immediately after sitting down?"
  • Multiple Choices with Plausible Distractors: Distractors (incorrect options) should be plausible and require watching the video carefully to rule out. Distractors may involve objects or events that appear similar to the correct answer or occur at different times in the video. For example: In the question "What does the person eat in the video?", possible choices could be:
    • A. A sandwich
    • B. An apple
    • C. A book (implausible distractor)
    • D. A banana (plausible distractor)
  • Cultural or Contextual Awareness: Ensure that questions are culturally sensitive and relevant to the video’s context, especially if the video pertains to specific cultural events or actions. This will help avoid bias and make the dataset more generalizable. For example: "What festival is being celebrated in the video based on the decorations and activities?"
  • MCQs Format: The MCQs should be written in the following format:
    • Question: <Question>
    • Answer: Correct_Answer (A. Option 1, B. Option 2, C. Option 3, D. Option 4)
  • Example MCQ:
    • Question: What traditional Japanese garment are the individuals in the Video wearing?
    • Answer: Kimono (A. Kimono, B. Sari, C. Hanbok, D. Cheongsam)
R.3 Instructions for Writing Short Answer Questions (SAQs)
  • Concise and Precise Questions: The questions should be brief and to the point, avoiding unnecessary complexity. For example: "What color is the car that drives by at the beginning?"
  • Culturally Relevant Questions: If the video depicts culturally specific events or actions, ensure that the question is contextually appropriate for that setting. This ensures relevance and avoids bias. For example: "What festival is being celebrated in the video?" (based on visual cues like decorations or attire).
  • Answer Length: Answers should be concise, generally between 1-10 words.
R.4 Verifying Phi-4 Scores
Volunteers are tasked with verifying the phi-4 scores assigned to model predictions for a set of videos. The verification process involves reviewing the video, the question, the ground truth answer, the model’s prediction, and the assigned score. Based on this review, you will determine whether the assigned score is appropriate or if it needs adjustment.
Step-by-Step Instructions
  1. 1.
    Watch the Video: Using the mint_video_id, watch the video associated with each QA.
  2. 2.
    Review the Question, Ground Truth Answer, and Model Prediction: Carefully read the question (Q), the ground truth answer (A), and the model’s prediction (Prediction) in the provided columns.
  3. 3.
    Check the Assigned Score: Look at the score already assigned in the Score column. The score ranges from 0 to 5, with increments of 0.5.
  4. 4.
    Assess the Model’s Prediction: Ask yourself: Does the model’s prediction deserve the score it was given compared to the ground truth answer?
  5. 5.
    Select Your Response:
    • If YES, indicating that the model’s prediction is appropriately scored, select YES from the dropdown menu under the Do you agree with the score? column. Leave the Your Score column empty.
    • If NO, indicating that the model’s prediction does not deserve the score assigned, select NO from the dropdown menu under the Do you agree with the score? column. Then, select your preferred score from the dropdown menu in the Your Score column.
R.5 Verifying Machine Translated Generic Category QA pairs
Volunteers are tasked with verifying the machine-translated QA pairs for different languages. There are two types of QA pairs: Open-ended and Multiple Choice Questions (MCQs). The goal is to verify the accuracy of the translations for both types of questions and answers. If the translation is correct, the volunteer will mark it as "YES"; if incorrect, they will correct the translation and provide the updated information in the appropriate columns.
Step-by-Step Instructions
  1. 1.
    Understand the Columns: The Excel sheet contains the following columns for English (Ground Truth):
    • English_Question
    • English_MCQ_Choice_1, English_MCQ_Choice_2, English_MCQ_Choice_3, English_MCQ_Choice_4
    • English_Answer
    The following columns are for the translations:
    • Translated_Question
    • Translated_MCQ_Choice_1, Translated_MCQ_Choice_2, Translated_MCQ_Choice_3, Translated_MCQ_Choice_4
    • Translated_Answer
  2. 2.
    Check the Translation Accuracy: The English text serves as the Ground Truth. For each row, review the translated versions of the question, MCQ choices, and answer.
  3. 3.
    Verify the Translation:
    • If the translation is correct, write YES in the Is the translation correct? column.
    • If the translation is incorrect, write NO in the Is the translation correct? column, and insert the correct translation in the respective columns:
      • Correct_Translated_Question
      • Correct_Translated_MCQ_Choice_1, Correct_Translated_MCQ_Choice_2, Correct_Translated_MCQ_Choice_3, Correct_Translated_MCQ_Choice_4
      • Correct_Translated_Answer
  4. 4.
    Handling Open-ended QA Pairs:
    • For Open-ended QA pairs, the MCQ columns will be empty.
    • If the question is Open-ended and the translation is incorrect, only fill in the Correct_Translated_Question column and leave the MCQ and answer columns empty.
Example
Incorrect Translation Example:
  • If the translated question is incorrect:
    • Write NO in the Is the translation correct? column.
    • Insert the correct translation in the Correct_Translated_Question column.
    • Leave the other columns (MCQ choices and answer) empty if they are correct.